2)Berman & Jaccor Corporation's current sales and partial balance sheet are shown below.

This year
Sales $ 1,000
Balance Sheet: Assets
Cash $ 150
Short-term investments $ 90
Accounts receivable $ 200
Inventories $ 100
    Total current assets $ 540
Net fixed assets $ 500
    Total assets $ 1,040

Sales are expected to grow by 12% next year. Assuming no change in operations from this year to next year, what are the projected total operating assets? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1
Answer: $1,064
Explanation and Calculation:
Total operating assets [current year] = Cash+Accounts receivable+Inventories+Net fixed assets = 150+200+100+500 = $                950
Assuming total operating assets also grow with sales, Projected total operating assets = 950*112% = $             1,064
[Operating assets do not include short term investments]